OpenStudy (anonymous):

Determine the solution of the system? vertically set up y=x+6 y=2x

OpenStudy (saifoo.khan):

2x = x + 6 Solve.

OpenStudy (saifoo.khan):

Substitute y = 2x in the first equation.

OpenStudy (anonymous):

2x=x+6

OpenStudy (albert0898):

Y = x + 6 When, Y = 2x [1] 2x = x + 6 [2] Subtract 'x' from both sides [3] x = 6